OpenStudy (anonymous):

Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved five thousand two hundred dollars to pay back to his grandfather. At what interest rate did Zach borrow the money? Round to the nearest whole percent. twelve percent fifteen percent eighteen percent thirty-eight percent

OpenStudy (anonymous):

Wait @heather040200 could you come back please

OpenStudy (anonymous):

yo

OpenStudy (anonymous):

lol okay so we have 1900=(4000x R x 3)/100 so would whats R?

OpenStudy (anonymous):

I = p r t <=> r = I/ p * t The difference amount between borrowed and paid back is the interest: I = $5,920 - $4,000 = $1,920 => Interest rate: r = 1,920 / 4,000 * 3 = .16 = 16%

OpenStudy (anonymous):

16 isnt an answer

OpenStudy (anonymous):

i rounded

OpenStudy (anonymous):

so would it be 15?

OpenStudy (anonymous):

ya